OpenStudy (anonymous):

Solve the solution: 5x+5-8(x+1)=6x+4 Now I do this and I get 0 but it's incorrect? I have no idea what I'm doing wrong, I've done this like 3 times and have gotten the same answer.

OpenStudy (anonymous):

-7/9

OpenStudy (anonymous):

5x + 5 -8x +8 = 6x +4 -3x -6x = 4 -8 -5 -9x = -9 ( - 1) x = 9/9 x = 1

OpenStudy (anonymous):

The answer is -7/9 but can you explain how you got that? @Jpenrod

OpenStudy (anonymous):

Nevermind, I see what I did wrong!

OpenStudy (anonymous):

ops! Just as I did, only wrong when multiplied -8 (x + 1)! my mistake -> -8 (x +1) =-8x + 8, the correct thing - 8x -8